question if the circumference of a cicle with r by c=2nr and n= 3.14 find the circumference when r = 15 meters, so far I have radius = 15 and n= 3.14 so r= 2*15 = 30 Im stuck....I just want to know if on right track.
Answer by stanbon(75887)

find the circumference when r = 15 meters,
----
c = 2(pi)r
c = 2(3.1416)*15
c = 30*3.1416
circumference = 94.248 meters
=================================
Cheers,
Stan H.
================